# Pralekha: Cross-Lingual Document Alignment for Indic Languages

### v2

@@ 6.1 Intrinsic Performance of DAC
DAC consistently achieves the highest precision across all settings, with gains of 15–20% over pooling-based baselines. While its recall is slightly lower, DAC attains competitive or superior F1 scores, indicating that it produces more precise alignments, a desirable property for high-quality MT training data, where quality is more important than quantity Ranathunga et al. (2024); Zhou et al. (2023). Among the baselines, LIDF is the strongest competitor, performing closest to DAC due to its effective combination of length normalization and term weighting. LP performs slightly worse, whereas MP and IDF-based pooling lag considerably behind. Additionally, in experiments conducted on the CCAligned dataset El-Kishky et al. (2020), DAC outperforms LIDF across all metrics; see Appendix B for more details.
@@
Effect of Granularity and Embedding Model:
Performance decreases slightly as $G$ increases, with the largest drop from $G=1$ to $G=2$; differences between $G=4$ and $G=8$ are minimal. This trade-off is expected, as higher granularities reduce the number of chunks to align, enabling faster mining at a small cost to quality. Among embedding models, SONAR performs better with pooling-based methods, whereas LaBSE achieves higher performance with DAC. This aligns with their training setups: SONAR is trained with mean-pooled embeddings, while LaBSE uses CLS representations, which are more effective for fine-grained alignment.
In conclusion, the optimal configuration for document alignment is LaBSE with DAC at granularities $G=1$, $G=2$, and $G=4$, followed closely by SONAR with LIDF at the same granularities. See Appendix A for a more detailed analysis of intrinsic performance trends.
Based on these findings, we use the selected methods for extrinsic evaluation.
@@ 6.2 Extrinsic Performance of DAC
Table 5 reports COMET scores for extrinsic evaluation, comparing DAC with LIDF, the strongest baseline from Section 6.1. DAC consistently outperforms LIDF across all granularities, translation directions, and fine-tuned models, showing that its higher-precision alignments produce better document pairs and improve translation quality. Detailed per-language COMET scores are also provided in Table 5 in the Appendix.
Differences in MT performance across granularities are small, consistent with intrinsic trends. The strong results for $G=2$ and $G=4$ indicate that chunk-based alignment can reduce computation while maintaining quality.
@@
Overall, performance differences are modest but consistent. Extrinsic results show that DAC-aligned documents at $G=2$ and $G=4$ achieve strong translation quality, while intrinsic performance does not always predict downstream gains.
Appendix C provides a more detailed analysis of extrinsic performance trends. Additionally, we report ChrF scores in Table 6, which follow trends similar to the COMET results.
@@ 7 Conclusion
Mining parallel document pairs for document-level MT is challenging because existing CLDA methods often rely on metadata or pooled sentence embeddings due to the limited context windows of embedding models. To address these challenges for Indic languages, we introduced Pralekha, a large-scale benchmark containing over 3 million aligned document pairs across 11 Indic languages and English. We also proposed the Document Alignment Coefficient (DAC), a fine-grained alignment metric that leverages chunk-level matches instead of traditional pooling-based representations. Intrinsic evaluation in realistic noisy settings shows that DAC outperforms existing baselines, achieving higher precision and F1 scores. Extrinsic evaluation further demonstrates that document-level MT models trained on DAC-aligned pairs consistently outperform those trained on baseline alignments, validating DAC’s effectiveness for parallel document mining. By releasing the Pralekha dataset and CLDA evaluation framework, we aim to enable further research on developing and evaluating document-level MT models for Indic languages.
@@
While Pralekha and the proposed Document Alignment Coefficient (DAC) advance research in cross-lingual document alignment (CLDA), several limitations remain. Our dataset covers 11 Indic languages and English, so the findings may not generalize to other language families with different linguistic properties, such as complex morphology or divergent word order. The documents in Pralekha are mainly drawn from structured, high-quality sources (PIB and MKB). As a result, the benchmark is less representative of noisy web data that contains informal language, OCR errors, or code-switching. In addition, the benchmark primarily includes documents of moderate length. Performance on much shorter or longer documents, which are common in other domains, remains untested. The DAC metric, though more accurate than pooling-based methods, is computationally expensive because it performs chunk-level alignment. This cost may limit its use for very large corpora or real-time applications. DAC may also be less robust when translations are semantically equivalent but highly paraphrased. Finally, our extrinsic evaluation does not fully establish how improvements in alignment quality affect document-level MT performance, since current MT metrics capture discourse-level phenomena only to a limited extent. Future work should expand CLDA evaluation to more diverse language families, noisier domains, and a wider range of document lengths. It should also explore alignment strategies that are more robust to paraphrasing, improve computational efficiency, and develop document-level MT metrics that better correlate with alignment quality.
@@
The code and datasets created in this work will be made available under permissible licenses. Generative AI systems were only used for assistance purely with the language of the paper, eg: paraphrasing, spell-check, polishing the author’s original content and for writing boiler plate code.
@@ Appendix B Intrinsic Evaluation on CCAligned
While the primary intrinsic evaluation was conducted on the Pralekha dataset, we also perform an additional benchmark on CCAligned El-Kishky et al. (2020), a large-scale web-mined parallel corpus. This experiment evaluates whether the performance gains of DAC generalize beyond the structured and high-quality domains of Pralekha. We follow the same data settings described in Section 5.1. For comparison, we use the strongest baseline from the main experiments: the LIDF pooling method with SONAR embeddings, evaluated at granularities $G=1,2,4$ and the DAC approach with LaBSE embeddings at the same granularities.

### v3

@@ 6.1 Intrinsic Performance of DAC
DAC consistently achieves the highest precision across all settings, with gains of 15–20% over pooling-based baselines. While its recall is slightly lower, DAC attains competitive or superior F1 scores, indicating that it produces more precise alignments, a desirable property for high-quality MT training data, where quality is more important than quantity ranathunga2024qualitydoesmatterdetailed; zhou2023limaalignment. Among the baselines, LIDF is the strongest competitor, performing closest to DAC due to its effective combination of length normalization and term weighting. LP performs slightly worse, whereas MP and IDF-based pooling lag considerably behind. Additionally, in experiments conducted on the CCAligned dataset el-kishky-etal-2020-ccaligned, DAC outperforms LIDF across all metrics; see Appendix B for more details.
Granularity LaBSE[Uncaptioned image] SONAR[Uncaptioned image]
Time (min) Precision Recall F1 Time (min) Precision Recall F1
G = 1 47 0.9152 0.6588 0.7635 115 0.9106 0.6484 0.7550
G = 2 36 0.9007 0.6411 0.7463 96 0.8870 0.6108 0.7210
G = 4 25 0.8936 0.6471 0.7480 59 0.8742 0.6005 0.7098
G = 8 18 0.8912 0.6407 0.7428 35 0.8732 0.6020 0.7109
Table 3: Computational time and alignment performance across document granularities ($G$) for 150k document pairs on an NVIDIA H100 node. Results for LaBSE and SONAR are reported in wall-clock time (minutes), precision, recall, and F1 score. Color gradients illustrate relative trends: redgreen indicates slower→faster configurations for time, while performance metrics follow a unified green scale. The drop in alignment quality is far smaller than the reduction in computation time, highlighting a strong computational advantage at higher granularities.
@@
Effect of Granularity.
As shown in Table 3, alignment performance is highest at the finest granularity ($G=1$) for both LaBSE and SONAR, with only a modest decline as chunk size increases. Beyond $G=2$, variations across performance metrics remain minimal, indicating that coarser segmentations preserve much of the alignment quality. In contrast, computational time decreases sharply from 47 to 18 minutes for LaBSE and from 115 to 35 minutes for SONAR as $G$ increases from 1 to 8, resulting in roughly a 2–3× reduction in runtime. This efficiency gain stems from a combinatorial reduction in the candidate search space: at higher granularities, each document is represented by fewer, longer chunks, leading to substantially fewer embedding computations and pairwise similarity operations during retrieval.
This trade-off underscores an important consideration for large-scale document alignment. While fine-grained ($G=1$) configurations yield slightly better alignment quality, their computational cost becomes impractical at scale. Coarser granularities thus offer a balanced solution, maintaining competitive alignment performance while enabling significantly faster processing, offering a favorable efficiency–performance trade-off .
Effect of Embedding Model.
Across embedding models, SONAR performs better with pooling-based methods, whereas LaBSE achieves higher performance with DAC. This contrast reflects their underlying training objectives: SONAR is trained with mean-pooled sentence representations, making it naturally compatible with aggregation-based alignment, while LaBSE relies on the [CLS]-token embedding, which encodes context from the entire sequence while remaining sensitive to token-level structure. Consequently, LaBSE benefits from methods that preserve token-level information, whereas SONAR aligns more effectively with pooling strategies that emphasize global semantics. See Appendix A for a detailed analysis of intrinsic performance trends.
@@ 6.2 Extrinsic Performance of DAC
Based on the intrinsic evaluations, the optimal configurations for document alignment on our Pralekha dataset are LaBSE with DAC at granularities $G=1$, $G=2$, and $G=4$, followed closely by SONAR with LIDF at the same granularities. These configurations were chosen solely on the basis of alignment performance metrics, without accounting for computational cost. In this section, we evaluate their extrinsic performance using the experimental setup described in Section 5.2, to examine whether the intrinsic alignment trends translate into measurable downstream gains.
Table 6 reports COMET scores for extrinsic evaluation, comparing DAC with LIDF (the strongest baseline). DAC consistently outperforms LIDF across all granularities, translation directions, and fine-tuned models, demonstrating that higher-precision alignments yield cleaner parallel data and result in measurable downstream improvements in translation quality. Detailed per-language COMET scores are presented in Appendix C.
Differences in MT performance across granularities are relatively small and align closely with intrinsic performance trends. The strong results for $G=2$ and $G=4$ suggest that moderate chunking can substantially reduce computational cost while maintaining translation quality. This suggests that coarse-grained segmentation balances alignment efficiency and performance in large-scale settings.
@@
Overall, performance differences remain modest but consistent across settings. Extrinsic results confirm that DAC-aligned corpora at $G$ = 2 & 4 achieve strong translation quality. However, it is worth noting that intrinsic performance does not always translate to downstream performance. Appendix C provides further discussion of these trends, and Table 7 reports corresponding ChrF scores, which follow patterns similar to the COMET results.
@@ 7 Conclusion
Mining parallel document pairs for document-level MT is challenging because existing CLDA methods often rely on metadata or pooled sentence embeddings due to the limited context windows of embedding models. Moreover, sentence-level alignment introduces a combinatorially large search space, leading to high computational costs that become impractical at scale. To address these challenges for Indic languages, we introduced Pralekha, a large-scale benchmark comprising over 3 million aligned document pairs across 11 Indic languages and English. Furthermore, we proposed the Document Alignment Coefficient (DAC), a novel metric for fine-grained document alignment. Unlike pooling-based methods, DAC aligns documents by matching smaller chunks and computes similarity as the ratio of aligned chunks to the average number of chunks in a pair. Intrinsic evaluation shows that our chunk-based approach is 2–3× faster while maintaining competitive performance, and that DAC achieves substantial gains over pooling-based baselines. Extrinsic evaluation further demonstrates that document-level MT models trained on DAC-aligned pairs consistently outperform those trained using baseline alignment methods. These results highlight DAC’s effectiveness for parallel document mining, balancing performance with computational efficiency. By releasing the Pralekha dataset and CLDA evaluation framework, we aim to facilitate future research on scalable and reliable document-level alignment and to advance the development of high-quality document-level MT systems for Indic languages.
@@
While Pralekha and the proposed chunk-based alignment approach advance research in cross-lingual document alignment, some limitations remain. Our dataset covers 11 Indic languages and English, so the findings may not generalize to other language families with different linguistic properties. The documents in Pralekha are primarily drawn from structured, high-quality sources, making the benchmark less representative of noisy web data that includes informal language, OCR errors, or code-switching. In addition, the benchmark mostly contains documents of moderate length, leaving performance on very short or long documents, common in other domains, largely untested. Because the dataset is domain-specific, MT models trained solely on it may overfit and exhibit reduced generalization, or even hallucinate when evaluated on out-of-domain data. DAC may also be less robust when translations are semantically equivalent but highly paraphrased. Finally, our extrinsic evaluation does not fully establish how improvements in alignment quality affect document-level MT performance, as current MT metrics capture discourse-level phenomena only to a limited extent. Future work should expand CLDA evaluation to more diverse language families, noisier domains, and a wider range of document lengths. It should also explore alignment strategies that are more robust to paraphrasing and develop document-level MT metrics that better correlate with alignment quality.
@@
All datasets used in this paper are available under permissible licenses, and we adhere strictly to their intended usage, maintaining full compliance with licensing requirements. Additionally, the code used for our evaluation framework will be made publicly available under the MIT License. [^11] https://opensource.org/licenses/MIT Generative AI systems were used solely to assist with the language of this paper, specifically for paraphrasing, spell-checking, polishing the authors’ original text, and generating boilerplate code.
Acknowledgements
We would like to thank EkStep Foundation and Nilekani Philanthropies for their generous grant towards building datasets, models, tools, and other resources for Indian languages. This work was partly supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists 25K21290.
@@ Appendix B Intrinsic Evaluation on CCAligned
While the primary intrinsic evaluation was conducted on the Pralekha dataset, we also perform an additional benchmark on CCAligned el-kishky-etal-2020-ccaligned, a large-scale web-mined parallel corpus. This experiment evaluates whether the performance gains of DAC generalize beyond the structured and high-quality domains of Pralekha. We follow the same data settings described in Section 5.1. For comparison, we use the strongest baseline from the main experiments: the LIDF pooling method with SONAR embeddings, evaluated at granularities $G=1,2,4$ and the DAC approach with LaBSE embeddings at the same granularities.